# Data Efficient Prediction of excited-state properties using Quantum Neural Networks

### v1

@@ Abstract
Understanding the properties of excited states of complex molecules is crucial for many chemical and physical processes. Calculating these properties is often significantly more resource-intensive than calculating their ground state counterparts. We present a quantum machine learning model that predicts excited-state properties from the molecular ground state for different geometric configurations. The model comprises a symmetry-invariant quantum neural network and a conventional neural network and is able to provide accurate predictions with only a few training data points. The proposed procedure is fully NISQ compatible. This is achieved by using a quantum circuit that requires a number of parameters linearly proportional to the number of molecular orbitals, along with a parameterized measurement observable, thereby reducing the number of necessary measurements. We benchmark the algorithm on three different molecules by evaluating its performance in predicting excited state transition energies and transition dipole moments. We show that, in many instances, the procedure is able to outperform various classical models that rely solely on classical features.
@@ I Introduction
To address these challenges we propose a quantum machine learning (QML) model that directly operates on the quantum mechanical ground state on a quantum computer (QC). The quantum state is measured using only commuting operators drastically reducing the overhead on the QC. The model includes a quantum neural network (QNN) and a classical neural network (NN) designed to efficiently calculate excited-state properties [^1] We use “excited-state property” not only for intrinsic properties of the excited state but also to excited state related properties such as transition energies. for different geometric configurations of a given molecule. The architecture of the QNN is constructed to explicitly account for the symmetries of the ground state. The proposed model is able to predict these properties with only a few training data points which avoids costly excited state simulations. The overall procedure is sketched in Figure 1.
@@
In a related work [19] the first transition energies and TDMs are predicted by performing measurements with different observables on the ground state after an non-parametrized entangling layer. The measurement values are used as input for an NN that outputs the target functions. In [20] a deep neural network determines the parameters of a QNN that prepares excited states using variational quantum deflation or the subspace search VQE. While these approaches share a similar rationale regarding the benefits of interpolation in this context, our model predicts excited-state properties with a significantly lower number of training data points, by directly learning on the molecular ground state.
The remainder of this work is organized as follows. In Section II, we motivate and discuss the construction of the model and its training procedure. In Section III, we demonstrate the efficacy of the model in predicting excited-state properties from grounds states of LiH, $\mathrm{H}_{2}$, and rectangular $\mathrm{H}_{4}$, and provide evidence for its data efficiency compared to several other models. Finally, in Section IV, we discuss our findings, the limitations of our model and study design, and propose directions for further research.
@@ II.1 Molecular States on a QC
We consider molecules in an isolated environment, resulting in a spin-degeneracy between the $\alpha$ and $\beta$ orbitals. Figure 2 illustrates three spin configurations and their corresponding Jordan-Wigner mapped states for the example of $\mathrm{H}_{2}$. Near the equilibrium the lowest energy configuration has the highest amplitude in the ground state. However, higher energy configurations also contribute to the ground state but with smaller amplitudes. Each configuration is associated with a computational basis state on the QC. The squares of the amplitudes of these computational basis states represent the probabilities of electrons occupying the corresponding spin-orbitals. As the molecule dissociates, the shapes of the orbitals change, leading to variations in the occupation probabilities of the spin orbitals. Consequently, the amplitudes of the computational basis states in the ground state change accordingly. In the following section, we design a QNN to read and process these amplitudes and their variations to derive the target value $y$ (in our case an excited-state property in a given geometric configuration) using training data points. For the theoretical justification for why it is possible to predict properties of excited states from the ground state, we refer to [19].
@@ II.2 Model Ansatz
An intuitive approach to derive excited-state properties from the ground state would be to use the QNN as a parametrized excitation operator and measure the resulting state with the observable corresponding to the desired property. However, there are two problems with this approach, which we address with specific design choices of our model.
First, the observable $O=\mathbf{w}\mathbf{P}$ [e.g., Eq. (3)] is generally a linear combination of many non-commuting Pauli-strings $P$ with coefficients $\mathbf{w}$. This leads to a significant overhead on the QC, as only commuting operators can be measured on the same quantum state. We address this issue by defining an observable that consists solely of commuting Pauli-strings, specifically strings that only contain Pauli-$Z$ operators. This observable has the form
$O(\mathbf{w})=w_{1}\mathds{1}+w_{2}(\mathrm{Z}_{0}+\mathrm{Z}_{n_{\mathrm{orb}}})+w_{3}\mathrm{Z}_{0}\mathrm{Z}_{n_{\mathrm{orb}}}+\ldots\,,$ (4)
where $\mathds{1}$ is the identity operator. The observable is constructed such that the result remains unchanged when the first $n_{\rm orb}$ qubits are exchanged with the last $n_{\rm orb}$ qubits, as discussed in Section II.1. In this work, the qubits corresponding to the first orbital contribute equally to the final sum. Additional measurements on the qubits corresponding to higher orbitals can be incorporated following this approach. Measuring $O(\mathbf{w})$ requires that the QNN learns a basis change where this simplified observable measures the corresponding target function. This approach drastically reduces the measurement-related overhead on the QC.
@@
The second challenge is that both the excitation operator and the observable generally depend on the geometric configuration of the given molecule. In this work, $R$ is a scalar describing the distance between the atoms of the molecule, but in general $R$ can be a multidimensional vector. Although one might hope that the $R$ dependency could be negligible or that the QNN could project into an $R$-independent space, we explicitly consider these $R$ dependencies by introducing a small classical NN. This NN takes $R$ as input and outputs the weights $w_{i}$ in Equation (4). The linear combination of these operators constitutes the output of the model. Therefore, the NN is designed to learn and account for the $R$ dependence of the excitation operator and the observable in the model. An example of the overall structure of the model is depicted in Figure 1.
For the QNN, we choose a structure that is designed to be efficiently trainable and to have a useful bias towards the overall operations it is meant to perform. To integrate the molecular symmetry into the QNN, we make the QNN invariant under the exchange of $\alpha$ and $\beta$-electrons. This can be achieved by letting all operators and observables, along with their parameters acting on the first $n_{\mathrm{orb}}$ qubits, also act on the last $n_{\mathrm{orb}}$ qubits.
@@
To enhance the trainability, the structure of the QNN is partly inspired by quantum convolutional neural networks, which are known to avoid barren plateaus and have a number of parameters that scale logarithmically [23]. Together with the given symmetry restriction the number of parameters is further reduced, scaling linearly with the number of orbitals considered. Specifically, when pooling to two qubits the number of parameters is $n_{\mathrm{params}}=8(n_{\mathrm{orb}}-1)-3$. The relevance of the findings in paper [24] regarding the simulatability of quantum convolutional neural networks is discussed in Chapter IV. The QNN consists of alternating two-qubit operation layers. The first layer (N) is designed for maximal expressivity, while the second layer (P) is a pooling layer designed to entangle the given qubits. After each pooling layer, the qubit with the larger index in the pooling operation is discarded. Qubits associated with orbitals of larger indices, as defined in Section II.1, are discarded earlier because the occupation probability in the near-equilibrium ground state generally decreases with increasing orbital indices. These operations are repeated until only two qubits remain, due to the small size of the molecules considered. However, for larger systems, one could repeat this process until the number of remaining qubits equals the number of electrons in the system. Our aim for the QNN is to transform the ground state into a state that, when measured, returns an approximation of the desired target state observable. Figure 3 shows an example of such a QNN for $\mathrm{H}_{2}$ with 4 orbitals. We call the resulting symmetry-invariant QNN combined with an NN “siQNN-NN”. Note that we have chosen the particular QNN architecture in Figure 3 for its computational efficiency. However, the proposed method of letting a QNN operate directly on ground states is not restricted to this specific choice.
@@
Figure 4: The target functions for LiH in (a) and (b), for $\mathrm{H}_{2}$ in (c) and (d), and for $\mathrm{H}_{4}$ in (e) and (f) in the 6-31G basis. In (a), (c), and (e) the transiton energies $\Delta E_{j}$ for the four energetically lowest excited singlet and triplet states, that do not exhibit avoided crossings, are shown in Hartree $\rm E_{\rm h}$. In (b), (d), and (f) the non-zero TDMs of those states are shown in units of elementary charge e and the Bohr radius $\rm{a}_{0}$. All values are obtained via exact diagonalization of Equation (3).
@@ II.3 Training
In the best case, the siQNN parameters are already optimal after pre-training, and the model output is similar to the target function. Then, the NN only fine-tunes the siQNN output towards the target function in regions where it is necessary to account for minor effects of the $R$ dependence. However, in some cases (e.g., the target function $\|\bm{\mu}{0,S_{2}}\|_{2}$ for LiH, see Figure 4) the siQNN is not able to adequately describe the target function on its own, due to its restricted expressiveness. Then, the pre-training might converge to parameters that are sub-optimal for the complete siQNN-NN model. In this case, the combined training becomes more important and the NN is used to account for the pronounced $R$-dependence.
In both training steps, we use an Adam [25] optimizer with the mean squared error (MSE) as a loss function. Training is terminated if one of three criteria is met: 1) if the training loss decreases below a predefined target loss $\mathcal{L}_{\rm t}$, 2) if the training loss does not improve over a certain number of iterations, or 3) if a maximum number of iterations is reached.
@@ III.1 Datasets
In selecting appropriate example molecules to benchmark the model, we try to satisfy a trade-off between sufficient complexity and reasonable computational costs. We therefore choose LiH which, with $n_{\mathrm{orb}}=5$, is small enough to be simulated, trained, and extensively studied while already exhibiting interesting electronic characteristics. Additionally, to analyze the scalability of our model we also benchmark it on $\mathrm{H}_{2}$ with $n_{\mathrm{orb}}=4$ and rectangular $\mathrm{H}_{4}$ with $n_{\mathrm{orb}}=6$.
@@
In our benchmarks, we consider two different excited-state properties for the targets $y$: transition energies and TDMs. The transition energies are given by $\Delta E_{j}=E_{j}-E_{S_{0}}$, where $j\in\{S_{1},S_{0},\dots,T_{1},T_{2},\dots\}$ labels the corresponding excited singlet ($S_{i}$) and triplet states ($T_{i}$). TDMs are given by
$\|\bm{\mu}{0,j}(R)\|_{2}=\|\langle\psi_{0}(R)|\bm{\mu}|\psi_{j}(R)\rangle\|_{2}\,.$ (5)
Here, $\|\cdot\|_{2}$ is the $L_{2}$ norm, |$\ket{\psi_{j}(R)}$ with $j\in\mathcal{S}$ is the $j$-th excited state, and $\bm{\mu}$ is the dipole moment operator. In calculating the TDMs, we account for possible sign changes to ensure differentiability across the chosen $R$ spectrum.
For each molecule, we calculate the first four excited singlet and triplet state energies and the corresponding non-zero TDMs, see Figure 4. The upper row displays the transition energies, while the lower row shows the TDMs. The functions exhibit different degrees of complexity. Thus, it can be expected that different numbers of data points are required to describe each function with similar accuracy. The TDMs for LiH are comparatively complex, whereas the transition energies for $\mathrm{H}_{4}$ and partially for $\mathrm{H}_{2}$ are relatively simple. Furthermore, several target functions for a given molecule mostly differ by a constant offset, and therefore the performance of the models on these functions are similar. To avoid learning problems, we only retain the states in which the entire $R$ region, from repulsion to dissociation, does not exhibit obvious avoided crossings [13]. During training, we scale $R$ and all target values to the interval $[-1,1]$.
In the following experiments, we assume that the molecular states on the QC are provided by some quantum algorithm. In the NISQ era, this could for example be done using VQE. To isolate the performance of the model from potential impurities in the input quantum states, we initialize the quantum register directly to state vectors obtained by diagonalizing the qubit Hamiltonian Equation (3).
We compare the performance of the siQNN-NN model to well-known classical models with a reduced feature space, i.e., $\overline{\mathcal{X}}=\mathcal{R}$. The models are trained only on this reduced feature space because the size of the state vector of a molecular ground state is $2^{2n_{\mathrm{orb}}}$. Even if the amplitudes were available classically (e.g., by state tomography), including the molecular states in $\mathcal{X}$ would make the feature dimension more than an order of magnitude larger than the size of the training dataset used in this study, which prevents a generalization for these models. However, in Section IV we argue, that it might be possible to construct a purely classical model, that manages to extract valuable information from the large state vector of the molecular states, but this would require a much more sophisticated ansatz and would therefore be an interesting follow-up research topic.

### v2

# thanks: Corresponding author

@@ Abstract
Understanding the properties of excited states of complex molecules is crucial for many chemical and physical processes. Calculating these properties is often significantly more resource-intensive than calculating their ground state counterparts. We present a quantum machine learning model that predicts excited-state properties from the molecular ground state for different geometric configurations. The model comprises a symmetry-invariant quantum neural network and a conventional neural network and is able to provide accurate predictions with only a few training data points. The proposed procedure is fully NISQ compatible. This is achieved by using a quantum circuit that requires a number of parameters linearly proportional to the number of molecular orbitals, along with a parameterized measurement observable, thereby reducing the number of necessary measurements. We benchmark the algorithm on three different molecules with three different system sizes: $H_{2}$ with four orbitals, LiH with five orbitals, and $H_{4}$ with six orbitals. For these molecules, we predict the excited state transition energies and transition dipole moments. We show that, in many cases, the procedure is able to outperform various classical models (support vector machines, Gaussian processes, and neural networks) that rely solely on classical features, by up to two orders of magnitude in the test mean squared error.
@@ I Introduction
To address these challenges we propose a quantum machine learning (QML)[19] model that directly operates on the quantum mechanical ground state on a quantum computer (QC). The quantum state is measured using only commuting operators drastically reducing the overhead on the QC. The model includes a quantum neural network (QNN) and a classical neural network (NN) designed to efficiently calculate excited-state properties [^1] We use “excited-state property” not only for intrinsic properties of the excited state but also to excited state related properties such as transition energies. for different geometric configurations of a given molecule. The architecture of the QNN is constructed to explicitly account for the symmetries of the ground state [20, 21, 22]. The proposed model is able to predict these properties with only a few training data points which avoids costly excited state simulations. The overall procedure is sketched in Figure 1.
@@
In a related work [23] the first transition energies and TDMs are predicted by performing measurements with different observables on the ground state after an non-parametrized entangling layer. The measurement values are used as input for an NN that outputs the target functions. In [24] a deep neural network determines the parameters of a QNN that prepares excited states using variational quantum deflation or the subspace search VQE. While these approaches share a similar rationale regarding the benefits of interpolation in this context, their models are trained with a significantly larger amount of training data (about a factor of 5 times the training data points our models uses) reaching similar accuracy as our model. Therefore, they do not compare themselves to classical interpolation methods, as we do.
The remainder of this work is organized as follows. In Section II, we motivate and discuss the construction of the model and its training procedure. In Section III, we demonstrate the efficacy of the model in predicting excited-state properties from the ground states of LiH, $\mathrm{H}_{2}$, and rectangular $\mathrm{H}_{4}$, and provide evidence for its data efficiency compared to several other models. Finally, in Section IV, we discuss our findings, the limitations of our model and study design, and propose directions for further research.
@@ II.1 Molecular States on a QC
We consider molecules in an isolated environment, resulting in a spin-degeneracy between the $\alpha$ and $\beta$ orbitals. Figure 2 illustrates three spin configurations and their corresponding Jordan-Wigner mapped states for the example of $\mathrm{H}_{2}$. Near the equilibrium the lowest energy configuration has the highest amplitude in the ground state. However, higher energy configurations also contribute to the ground state but with smaller amplitudes. Each configuration is associated with a computational basis state on the QC. The squares of the amplitudes of these computational basis states represent the probabilities of electrons occupying the corresponding spin-orbitals. As the molecule dissociates, the shapes of the orbitals change, leading to variations in the occupation probabilities of the spin orbitals. Consequently, the amplitudes of the computational basis states in the ground state change accordingly. In the following section, we design a QNN to read and process these amplitudes and their variations to derive the target value $y$ (in our case an excited-state property in a given geometric configuration) using training data points. For additional theoretical justification for why it is possible to predict properties of excited states from the ground state, we refer to [23].
@@ II.2 Model Ansatz
The intuitive approach to derive the transition energy $\Delta E$ to an excited state from the ground state would be to find the unitary excitation operator $U_{j}^{\rm exc}$ that transitions the ground state to the $j$-th excited state and measuring the Hamiltonian $H$ of the system in the excited state. From the result, one can subtract the measurement of $H$ in the ground state to obtain $\Delta E$. However, there are three problems with this approach that we address with the specific design choices of our model.
Firstly, the structure of $U_{j}^{\rm exc}$ is generally not known, and an exact derivation would introduce a significant computational overhead. Therefore, we aim to model its action with a machine learning approach consisting of a QNN with $U_{\rm QNN}$ acting on the ground state.
Secondly, in general, a large number of shots are required to derive $\Delta E=E_{j}-E_{0}$ by measuring $E_{0}$ and $E_{j}$ before subtracting. This is because for larger molecules $\Delta E$ can be on a lower energy scale than $E_{0}$ and $E_{j}$ and if one wants to know $\Delta E$ to a certain accuracy, $E_{0}$ and $E_{j}$ need to be derived to a larger accuracy, which leads to a larger number of shots required, compared to a direct measurement of $\Delta E$. Therefore, we aim to directly approximate the transition energy Hamiltonian $\tilde{H}$, i.e.,
$\Delta E$ $=\langle\psi_{0}|\left(U_{j}^{\rm exc}\right)^{\dagger}HU_{j}^{\rm exc}|\psi_{0}\rangle-\langle\psi_{0}|H|\psi_{0}\rangle$
$=\langle\psi_{0}|\left(U_{j}^{\rm exc}\right)^{\dagger}HU_{j}^{\rm exc}-H|\psi_{0}\rangle\equiv\langle\psi_{0}|\tilde{H}|\psi_{0}\rangle\,,$ (4)
where $U_{\mathrm{diag}}$ is the unitary that diagonalizes $\tilde{H}$, so that we can measure its eigenvalues in the computational basis with the measurement $D$. Since the Hamiltonian can be expressed as a linear combination of many non-commuting Pauli-strings [see Equation (3)] the same holds true for $\tilde{H}$ in Equation (4), because $U_{j}^{\rm exc\dagger}HU_{j}^{\rm exc}$ is also a linear combination of (potentially different) Pauli-strings. Theoretically, it is possible to derive this unitary operation $U_{\mathrm{diag}}$ that diagonalizes $\tilde{H}$, so that we can measure its eigenvalues (the $\Delta E$) in the computational basis with
$O(\mathbf{w})=w_{1}\mathds{1}+w_{2}(Z_{0}+Z_{n_{\mathrm{orb}}+1})+w_{3}Z_{1}Z_{n_{\mathrm{orb}}+1}+\ldots\,,$ (5)
where $\mathds{1}$ is the identity. Note that this observable explicitly respects the spin symmetry discussed in Section II.1. Since we do not know the exact structure of $\tilde{H}$ we leave it to the QNN to rotate into its eigenbasis and to derive the weights $w_{i}$ of the resulting measurements in this basis. We thus approximate $\Delta E$ by measuring
$\Delta E\approx\bra{\psi_{0}}U_{\rm QNN}^{\dagger}O(\mathbf{w})U_{\rm QNN}\ket{\psi_{0}}\,.$ (6)
Depending on the complexity of the problem, we might not find the exact eigenbasis of the full $\tilde{H}$ but only its leading contributions.
@@
The third challenge is that the coefficients in $H$ and therefore also in $\tilde{H}$ generally depend on the geometric configuration $R$ of the given molecule. In this work, $R$ is a scalar describing the distance between the atoms of the molecule, but in general $R$ can be a multidimensional vector. Although one might hope that the $R$ dependency could be negligible or that the QNN could project into an $R$-independent space, we explicitly consider this $R$ dependence by introducing a small classical NN. This NN takes $R$ as input and outputs the weights $w_{i}$ in Equation (5). The linear combination of these operators constitutes the output of the model. Therefore, the NN is designed to learn and account for the dependence of $R$ in $\tilde{H}$. An example of the overall structure of the model is depicted in Figure 1. All the same arguments that lead to Equation (6) can be used to derive general matrix elements connecting the ground state to the $j$-th excited state of an operator of interest $V$, i.e.,
$\|V_{i,j}\|_{2}$ $=\|\langle\psi_{j}|V|\psi_{i}\rangle\|_{2}=\sqrt{\langle\psi_{i}|V^{\dagger}|\psi_{j}\rangle\langle\psi_{j}|V|\psi_{i}\rangle}$ (7)
$\equiv\langle\psi_{i}|\tilde{V}|\psi_{i}\rangle\,,$ (8)
where we have introduced $\tilde{V}=\sqrt{V^{\dagger}|\psi_{j}\rangle\langle\psi_{j}|V}$ in the last equation. In this work, we are primarily interested in obtaining the $L_{2}$-norm of the matrix element which is denoted by $\|\cdot|\|_{2}$ in the equation above. Note that other expectation values and matrix elements can be approximated similarly. For the structure of the QNN we aim for expressiveness, so that the QNN can implement the necessary operations, and efficient trainibility with a useful bias towards the overall operations it is meant to perform, so that the correct operations are chosen. To integrate the molecular symmetry into the QNN, we consider only spin-symmetric operations. This is achieved by using the same operations on the first $n_{\mathrm{orb}}$ and last $n_{\mathrm{orb}}$ qubits.
@@
To enhance trainability, the structure of the QNN is partially inspired by quantum convolutional neural networks, which are known to avoid barren plateaus and have a circuit depth that scales logarithmically [27]. Together with the given symmetry restriction, the number of parameters is further reduced, scaling linearly with the number of orbitals considered. Specifically, when pooling to two qubits, the number of parameters is $n_{\mathrm{params}}=8(n_{\mathrm{orb}}-1)-3$ (see Appendix A). The QNN consists of alternating two-qubit operation layers. The first layer (N) is designed for maximal expressivity between neighboring qubits and is intended to implement rotations between adjacent orbitals inspired by ADAPT-VQE structural approaches (compare, e.g. [28]), while the second layer (P) is a pooling layer designed to entangle the given qubits. After each pooling layer, half the qubits are discarded and the information is pooled onto the remaining qubits. These operational layers are repeated until only a small number of qubits remain. Note that for large systems, this process could be repeated until the number of remaining qubits equals the number of electrons in the system. In this case, additional contributions in Equation 5 are further Pauli-strings consisting of $Z$-operators, respecting the given symmetry, and acting on qubits which are left in the circuit after the pooling operations (indicated by the dots in the equation). Figure 3 shows an example of such a QNN for $\mathrm{H}_{2}$ with four orbitals. We call the resulting symmetry-invariant QNN combined with an NN “siQNN-NN”. Quantum convolutional neural networks are known to be simulatable classically [29]. The relevance of these findings, namely the fact that this is not trivially possible but might be an interesting further research direction, is discussed in Section IV. Note that we have chosen the particular QNN architecture in Figure 3 for its computational efficiency. However, the proposed method of letting a QNN operate directly on ground states is not restricted to this specific choice.
@@
Figure 4: The target functions for LiH in (a) and (b), for $\mathrm{H}_{2}$ in (c) and (d), and for $\mathrm{H}_{4}$ in (e) and (f) in the 6-31G basis. In (a), (c), and (e) the transition energies $\Delta E_{j}$ for the four energetically lowest excited singlet and triplet states, that do not exhibit avoided crossings, are shown in Hartree $\rm E_{\rm h}$. In (b), (d), and (f) the non-zero TDMs of those states are shown in units of elementary charge e and the Bohr radius $\rm{a}_{0}$. All values are obtained via exact diagonalization of Equation (3).
@@ II.3 Training
In the best case, the siQNN parameters are already optimal after pre-training, and the model output is similar to the target function, because a good approximation to the eigenbasis of $\tilde{H}$ is found. Then, the NN only fine-tunes the siQNN output towards the target function in regions where it is necessary to account for minor effects of the $R$ dependence and potentially small contributions from a small term in $\tilde{H}$. However, in some cases (for example, the target function $\|\bm{\mu}{0,S_{2}}\|_{2}$ for LiH, see Figure 4) the siQNN is unable to adequately describe the target function on its own, due to its restricted expressiveness. Then, the pre-training might converge to parameters that are sub-optimal for the complete siQNN-NN model. In this case, the combined training becomes more important and the NN is used to account for the pronounced $R$ dependence and additional term contributions in $\tilde{H}$.
In both training steps, we use an Adam [30] optimizer with mean squared error (MSE) as a loss function. Training is terminated if one of three criteria is met: 1) if the training loss decreases below a predefined target loss $\mathcal{L}_{\rm t}$, 2) if the training loss does not improve over a certain number of iterations, or 3) if a maximum number of iterations is reached.
@@ III.1 Datasets
In selecting appropriate example molecules to benchmark the model, we try to satisfy a trade-off between sufficient complexity and reasonable computational costs. We therefore choose LiH which, with $n_{\mathrm{orb}}=5$, is small enough to be simulated, trained, and extensively studied while already exhibiting interesting electronic characteristics. Furthermore, to analyze the scalability of our model, we also benchmark it on $\mathrm{H}_{2}$ with $n_{\mathrm{orb}}=4$ and rectangular $\mathrm{H}_{4}$ with $n_{\mathrm{orb}}=6$.
@@
In our benchmarks, we consider two different excited-state properties for the targets $y$: transition energies and TDMs. The transition energies are given by $\Delta E_{j}=E_{j}-E_{S_{0}}$, where $j\in\{S_{1},S_{0},\dots,T_{1},T_{2},\dots\}$ labels the corresponding excited singlet ($S_{i}$) and triplet states ($T_{i}$). For the TDMs, we calculate the $L_{2}$ norm of the matrix element connecting the ground state to the $j$-th excited state. This corresponds to choosing $V=\bm{\mu}$ in Equation (8), where $\bm{\mu}$ is the dipole moment operator. In calculating the TDMs, we take into account possible sign changes to ensure differentiability across the chosen $R$ spectrum.
For each molecule, we calculate the first four excited singlet and triplet state energies and the corresponding non-zero TDMs, see Figure 4. The upper row displays the transition energies, while the lower row shows the TDMs. The functions exhibit different degrees of complexity. Thus, it can be expected that different numbers of data points are required to describe each function with similar accuracy. The TDMs for LiH are comparatively complex, whereas the transition energies for $\mathrm{H}_{4}$ and partially for $\mathrm{H}_{2}$ are relatively simple. Furthermore, several target functions for a given molecule differ mainly by a constant offset, and therefore the performance of the models on these functions is similar. To avoid learning problems, we only retain the states in which the entire $R$ region, from repulsion to dissociation, does not exhibit obvious avoided crossings [13]. During training, we scale $R$ and all target values to the interval $[-1,1]$.
In the following experiments, we assume that the molecular states on the QC are provided by some quantum algorithm. In the NISQ era, this could, for example, be done using VQE. To isolate the performance of the model from potential impurities in the input quantum states, we initialize the quantum register directly to state vectors obtained by diagonalizing the qubit Hamiltonian Equation (3).
We compare the performance of the siQNN-NN model to well-known classical models with a reduced feature space, i.e., $\overline{\mathcal{X}}=\mathcal{R}$. The models are trained only on this reduced feature space because the size of the state vector of a molecular ground state is $2^{2n_{\mathrm{orb}}}$. Even if the amplitudes were available classically (e.g., by state tomography), including the molecular states in $\mathcal{X}$ would make the feature dimension more than an order of magnitude larger than the size of the training dataset used in this study, preventing generalization for these models. However, in Section IV we argue that it might be possible to construct a purely classical model that manages to extract valuable information from the large state vector of the molecular states, but this would require a much more sophisticated ansatz and would therefore be an interesting follow-up research topic.
@@
III.2 Model Details
For the siQNN-NN, we use an NN with a single hidden layer comprising 41 parameters with a “tanh” activation function for the hidden layer. The NN is implemented using PyTorch [33], while the QNN is implemented in sQUlearn [34] and Qiskit, and then integrated into the model and trained using PennyLane [35] with state vector simulations. In Appendix B, we additionally show calculations for simulations including shot noise and find promising results. Pre-training is performed with a large learning rate of 0.5 and a maximum of 1,000 iterations to broadly explore the optimization landscape and prevent getting stuck in local minima. The subsequent training of the siQNN-NN is conducted with a learning rate between 0.005 and 0.05 and a maximum of 2,000 iterations to enable smooth convergence. The performance of the model learning on the ground state derived after pre-training labeled siQNN is also compared to the performance of the classical models in the following.
We compare the performance of these models to a purely classical NN that has the same structure as the siQNN-NN, but without the quantum component. The model is trained on the reduced feature space $\overline{\mathcal{X}}$. To ensure comparability, the number of parameters in the NN is equal to the total number of parameters in the siQNN-NN (the sum of parameters in the QNN and the NN). Choosing the same architecture as in the siQNN-NN allows us to isolate the contributions of the classical NN and QNN to make sure that the performance of the siQNN-NN does not solely rely on the classical part of the model. This model is trained with an ADAM optimizer until convergence (if over 1,000 iterations with 5 different learning rates between 0.01 and 0.1 the target loss is not improving) or until the target training loss $\mathcal{L}_{\rm t}$ is reached. Additionally, a support vector regressor (SVR) and a Gaussian process regressor (GPR) with a radial basis function kernel are implemented using scikit-learn [36]. Their hyperparameters are optimized using a grid search, where the performance of the hyperparameters is evaluated on the training dataset.
[…]